OpenStudy (anonymous):

Web 2.0 Tools User: Lewis Poole In Course: M/J Mathematics 2 V9 ( 2820) Instructor: Mrs. Kendra Nassar [E-Mail] You have 1 Unread E-Mail Message 06.05 Converting Between Measurement Systems Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Question 1 (Essay Worth 4 points) Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature.

OpenStudy (anonymous):

that moment when we cant help you on exams

OpenStudy (anonymous):

@Stitch_ might not be able to, but I can.

OpenStudy (anonymous):

The conversion formula is C = (5/9)*(F-32); where C = temperature in celcius and F = Temperature in Fahrenheit. So if we fill it in it looks like this: (0) = (5/9)*(32-32) Then we work it out. (0) = (5/9)*0 0 = 0 And there's your answer 32 degrees Fahrenheit equals zero degrees Celsius.

OpenStudy (anonymous):

You can give medals by clicking 'Best Response'.